# Calibration in Deep Learning: A Survey of the State-of-the-Art

### v3

@@ 4.2.3 Implicit Regularization: Differentiable Calibration Proxy
Recall that the aforementioned methods use a penalty term (either explicitly or implicitly) to improve model calibration on dataset $D$. There is a rising direction that directly optimizes objective function by using calibration errors (CE) as differentiable proxy  (?, ?) to standard loss:
@@
The focus of this line of work is to find differentiable approximations to calibration errors.   Kumar et al.  (?) proposed a kernel-based approach to explicitly calibrate models in training phrase called Maximum Mean Calibration Error (MMCE), which is differentiable and can be optimized using batch stochastic gradient algorithms. They cast the calibration error to be differentiable by defining an integral probability measure over functions from a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ induced by a universal kernel $k(\cdot,\cdot)$ and cannonical feature map $\phi:[0,1]\rightarrow\mathcal{H}$:
@@
Mukhoti and Gal (?) proposed patch accuracy versus patch uncertainty (PAvPU) for evaluate uncertain estimates for each image on segmentation task; Motivated by this idea,  (?) extended it to a differentiable proxy, accuracy versus uncertainty calibration (AvUC), which approximations to the accuracy versus uncertainty (AvU):
@@
Similarly (?) presented MDCA (Multi-class Difference in Confidence and Accuracy (MDCA)) for improving model calibration. It is a differentiable calibration loss function and independent of the binning mechanism. The idea was inspired by Static Calibration Error (SCE) (?). The authors showed the the combination with focal loss lead the least calibration error.
Remarks: Regularization methods, as compared to post-hoc methods, can directly output a well-calibrated model without additional steps. The increased complexity is different for different methods; for instance, focal loss can be seen as an implicit regularization and does not introduce observable additional computational complexity. Kernel-based and meta-network regularization add additional computations depending on the designed kernel methods and meta-networks.
@@ 4.2.4 Implicit Regularization: Data Augmentation
Remarks: This line of work combats overfitting by data augmentation in hidden space. This improves not only model generalization but also calibration. Those methods don’t significantly increase network complexity but usually require more training time due to more generated or synthesized training samples.
@@ 4.3 Uncertainty Estimation
This line of work aims to alleviate model miscalibration by injecting randomness.The popular methods are (1) Bayesian neural networks (?, ?), (2) ensembles (?), (3) Monte Carlo(MC) dropout (?) and (4) Gumbel-softmax (?) based approaches (?, ?). The former three sub-categories have been discussed in recent surveys (?, ?). Figure 4 presents the difference of those methods.
Figure 4: The methods of uncertainty estimation (?). (a) Bayesian neural network; (b) MC dropout; (c) Ensembles; (d) Gumbel-Softmax trick.
@@ 4.3.1 Bayesian Neural Networks
[…]
@@
4.3.2 MC Dropout, Ensembles and Gumbel-Softmax
By following the above-mentioned strategy, MC-dropout (?), ensembles (?) and Gumbel-softmax sampling (?, ?) introduce randomness in different ways. Then the $T$ in equation (30) corresponds to the number of sets of mask vectors from Bernoulli distribution $\{r^{t}\}_{t=1}^{T}$ in MC-dropout, or the number of randomly trained models in Ensembles, which potentially leads to different sets of learned parameters $\omega=\{\omega_{1},...,\omega_{t}\}$, or the number of sets of sampled attention distribution from Gumbel distribution $\{g^{t}\}_{t=1}^{T}$.
Remarks: This line of work requires multiple inference runs to perform approximations. This increases computational overhead significantly as compared to previous methods. On the other hand, besides model calibration, those methods are primarily proposed for uncertainty quantification and estimation. Therefore, network uncertainty can be captured and measured as well.
@@ 4.4 Hybrid Calibration Methods
Beside applying each method independently, we can always have calibration compositions by combining two or more methods. One straightforward way to combine non-post-hoc methods with post-hoc methods. For instance, performing Temperature Scaling (TS) after employing the regularization method and implicit calibration (?, ?). (?) observed that the combination of label smoothing and mixup training significantly improved calibration. While there are several possibilities for combining different approaches, we highlight some interesting compositions.
@@ 4.4.1 Ensemble Temperature Scaling (ETS)
(?) gave three important definitions related to calibration properties: accuracy-preserving, data-efficient, and expressive. They pointed out that TS is an accuracy-preserving and data-efficient approach but is less expressive. Ensemble Temperature Scaling (ETS) was proposed to improve TS expressivity while maintaining the other two properties:
@@ 4.4.2 Temperature Scaling with MC Dropout
(?) extended TS to dropout variational inference to calibrate model uncertainty. The key idea is to insert $\tau$ before final softmax activation and insert TS with $\tau>0$ before softmax activation in MC integration:$\hat{p}=\frac{1}{N}\sum_{i=1}^{N}\textsc{softmax}(\frac{f_{w_{i}}(x)}{\tau})$ where $N$ forward passes are performed to optimize $\tau$ with respect to NLL on the validation set. Then the entropy of the softmax likelihood is used to represent the uncertainty of all $C$ classes.
@@
Remarks: Appropriately combining different calibration types to some degree can further improve calibration, but, it may also combine their disadvantages, for instance, Ensemble Temperature Scaling (ETS) (?) has increased complexity.
@@ 4.5.1 Summary of Strengths and Limitations
(1) Implementation and Tuning Complexity. Implementing regularization method is less straightforward as compared to post-hoc ones. Also this line of work usually introduce hyper-parameters, for example, smoothing factor in label smoothing (?) and  (?) empirically showed that the importance of $\gamma$ in calibrating deep models. However, selecting the appropriate hyper-parameters can be challenging and may require extensive hyper-parameter tuning.
Uncertainty Estimation (UE) (1) Uncertainty quantification. Bayesian-method based calibration can quantify the uncertainty in the calibrated model (?). (2) Incorporating prior information. UE methods allow incorporation of prior knowledge or beliefs about the parameters being estimated and provide more robust estimates, especially when dealing with limited data (?). (1) Computational complexity. Many uncertainty estimation methods involve computationally intensive procedures, which can be impractical for large datasets or real-time applications, limiting their scalability. For example MC-dropout (?) requires multiple inference runs and perform averaging on predictions.
Hybrid Methods (1) Combines strengths: Hybrid methods can leverage the advantages of multiple approaches, potentially leading to better overall performance by addressing the limitations of individual techniques. Section 4.4 describes some of these methods (1) Increased complexity: Hybrid methods usually introduce additional complexity in model calibration, Integrating multiple techniques into a cohesive framework may pose implementation challenges and increase the risk of errors requiring more extensive validation and testing to ensure robustness and effectiveness.
@@ 4.5.2 Backbone Models and Datasets
Table 2: The state-of-the-art calibration methods and their categorization, datasets and backbone models to be calibrated with those methods.
4.6 Calibrating Pre-trained Large Models
@@
4.6.1 Large Vision Model Calibration
(?) studied recent state-of-the-art vision models that include vision transformer (?) and MLP-mixer (?). They found out that the model size and amount of pre-training in the recent model generation could not fully explain the observed decay of calibration with distribution shift or model size in the prior model generation. They also discussed the correlation between in-distribution and out-of-distribution (OOD) calibration. They pointed out that the models with better in-distribution calibration also gave better calibration on OOD benchmarks. (?) studied the calibration of CLIP (?) as a zero-shot inference model and found that CLIP is miscalibrated in zero-shot settings.They showed effectiveness of learning a temperature on an auxiliary task and applying it to inference regardless of prompt or datasets.
4.6.2 Large Language Models (LLMs)
The LLMs and prompting engineering have become an efficient learning paradigm and can perform numerous natural language tasks without or with only few examples. However, the outcome of this learning paradigm can be unstable and introduces bias with various prompt templates and training examples (?). The introduced biases include majority label bias, recency bias and token bias. To mitigate this bias, (?) proposed contextual calibration procedure to improve the predictive power of GPT-3 on few-shot learning tasks with a context-free input such as ”N/A”. The contextual calibration is performed by using vector scaling (?).
@@ 4.6.2 Large Language Models (LLMs)
Later on  (?) proposed domain-context calibration (DC) to mitigate label biases in in-context learning (ICL). The method tries to use random words sampled from the task corpus to estimate the effects of different label biases:
@@
where $W$ is a diagonal matrix. Similarly, (?) investigated the selection bias of LLMs in multi-choice question (MCQ) task and pinpointed that LLMs’ token bias is an intrinsic cause of the selection bias. They proposed label-free, inference-time debasing method PriDe, which demonstrates notable effectiveness and efficiency, interpretability, and cross-domain transferability. (?) extended mixup (?) training to improve model calibration by synthesizing samples based on the Area Under the Margin (AUM) for pre-trained language models.
Prototypical calibration (PROCA)  (?) is one of the latest studies in calibrating LLMs. It showed the importance of decision boundary in few-shot classification settings and suggested learning a better boundary with a prototypical cluster. Concretely, it estimates $K$ category-wise clusters with the Gaussian mixture model (GMM):
$P_{GMM}(\hat{p})=\sum_{k=1}^{K}\alpha_{k}p_{G}(\hat{p}|u_{k},\Sigma_{k})$ (36)
$p_{calibrated}=\frac{1}{T}\sum_{t=1}^{T}P_{GMM}(\hat{p}_{t})$ (37)
where $u_{k}$ and $\Sigma_{k}$ are the mean vector and covariance matrix of the distribution. The parameters $\{\alpha,u,\Sigma\}_{k=1}^{K}$ are estimated by using the Expectation-Maximization (EM) algorithm (?) with a small unlabelled dataset.
Most recently, (?) identified the limitations of previous approaches, for example, contextual and domain-context calibration have limitations in multi-sentence classification when the estimation of the contextual bias is inaccurate, and prototypical calibration is vulnerable to overfitting problem. Batch Calibration was then proposed to estimate the contextual bias for each class.
4.7 Calibration Applications
@@
4.7.1 Medical Imaging and Analysis
Calibrated deep learning models play a crucial role in medical imaging tasks such as disease diagnosis and medical image analysis (?). These models provide accurate predictions and uncertainty estimates, assisting clinicians in making informed decisions about patient care. For instance, calibrated models improve the reliability of automated diagnoses in radiology, pathology, and other medical imaging domains. By aligning predicted probabilities with observed outcomes, these models enhance the trustworthiness of automated medical decision support systems. Numerous studies have demonstrated the effectiveness of model calibration techniques in improving the accuracy and interpretability of medical imaging models.  (?) presents a network calibration procedure to make be robust to label noise based on observation that the confusion matrix of the noisy labels can be expressed as the matrix product between the confusion matrix of the clean labels and the label noises. (?) identified the issue of discrimination and calibration biases in models trained for automatic detection of malignant dermatological conditions from skin lesions images. They found that calibration metrics may appear to show that model outputs do not fairly represent uncertainty.
4.7.2 Financial Decision Making

### v4

@@ 4.2.3 Implicit Regularization: Differentiable Calibration Proxy
Recall that the aforementioned methods use a penalty term (either explicitly or implicitly) to improve model calibration on dataset $D$. There is a rising direction that directly optimizes objective function by using calibration errors as differentiable proxy  (?, ?) to standard loss:
@@
The focus of this line of work is to find differentiable approximations to calibration errors.
Kumar et al.  (?) proposed a kernel-based approach to explicitly calibrate models in training phrase known asMaximum Mean Calibration Error (MMCE), which is differentiable and can be optimized using batch stochastic gradient algorithms. They cast the calibration error to be differentiable by defining an integral probability measure over functions from a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$ induced by a universal kernel $k(\cdot,\cdot)$ and cannonical feature map $\phi:[0,1]\rightarrow\mathcal{H}$:
@@
Mukhoti and Gal et al. (?) proposed patch accuracy versus patch uncertainty (PAvPU) for evaluating uncertain estimates for each image on segmentation task; Motivated by this idea,  (?) extended it to a differentiable proxy, accuracy versus uncertainty calibration (AvUC), which approximates to the accuracy versus uncertainty (AvU):
@@
Similarly, Multi-class Difference in Confidence and Accuracy (MDCA) (?) was proposed to improve model calibration. MDCA is a differentiable calibration loss that does not rely on binning mechanisms, and its formulation is inspired by Static Calibration Error (SCE) (?). The authors demonstrated that combining MDCA with focal loss results in the lowest calibration error among the evaluated methods.
@@ 4.2.4 Implicit Regularization: Data Augmentation
4.2.5 Practical Considerations
Regularization-based calibration methods differ from post-hoc approaches by directly integrating calibration objectives into the training process, enabling models to produce well-calibrated predictions without separate correction steps (?, ?). The computational complexity varies considerably: simple implicit regularizers like focal loss (?) introduce negligible overhead, making them practical for large-scale training. In contrast, kernel-based methods such as Maximum Mean Calibration Error (MMCE) (?) and meta-learning approaches like differentiable ECE (DECE) (?) require additional computations related to kernel evaluations or meta-network optimization, which can increase training time.
Differentiable calibration proxies, including AvUC (?) and MDCA (?), offer flexible and efficient alternatives to traditional binning-based metrics by enabling end-to-end optimization of calibration errors. These methods have shown promising empirical improvements but can depend heavily on careful design and tuning to avoid overfitting (?). Overall, regularization methods provide a principled framework for calibration-aware training but may trade off simplicity and efficiency for potentially better calibration quality.
Implicit regularization via data augmentation improves calibration by enhancing generalization and reducing overfitting without modifying model architecture. Techniques like label smoothing (?) and mixup (?) are lightweight and easy to implement, and when combined, they often yield stronger calibration performance (?). However, they introduce additional hyperparameter (e.g., smoothing and mixing coefficients) that require careful tuning. These methods may also increase training time due to synthetic sample generation and can occasionally lead to underfitting if over-applied. Despite these trade-offs, they offer a practical, low-cost means to improve model calibration alongside robustness.
@@ 4.3 Uncertainty Estimation
This line of work aims to alleviate model miscalibration by injecting randomness.The popular methods are (1) Bayesian neural networks (?, ?), (2) ensembles (?), (3) Monte Carlo (MC) dropout (?) and (4) Gumbel-softmax (?) based approaches (?, ?). The former three sub-categories have been discussed in recent surveys (?, ?).
@@ 4.3.1 Bayesian Neural Networks
$\mathcal{L_{\theta}}=\int q_{\theta}(w)p(Y|X,w)dw-\textsc{KL}[q_{\theta}(w)\parallel p(w)]$ (32)
@@
4.3.2 Stochastic and Gradient-Based Uncertainty Methods
Several uncertainty estimation techniques improve calibration by introducing stochasticity or reweighting gradients during training. MC Dropout (?), deep ensembles (?, ?), and Gumbel-softmax sampling (?, ?) inject randomness in different ways: via dropout masks, independently trained model weights, or sampled attention distributions, respectively. These methods typically rely on multiple forward passes ($T$ samples) to approximate predictive uncertainty.
More recently, Lin et al. (?) proposed a sample-weighted gradient approach, BSCE-GRA, which scales gradients based on Brier Scores across all logits. Unlike traditional loss weighting, this method directly adjusts gradient magnitudes to prioritize uncertain samples without hindering convergence. Empirical results show that BSCE-GRA achieves competitive calibration while preserving model accuracy.
4.3.3 Practical Considerations
Uncertainty estimation methods enhance calibration by modeling predictive confidence through stochastic techniques like Bayesian neural networks (?), MC Dropout (?), ensembles (?), and Gumbel-softmax sampling (?). These approaches capture both aleatoric and epistemic uncertainty and improve reliability in high-risk settings. Gradient-based methods such as BSCE-GRA (?) further refine training by prioritizing uncertain samples without modifying model architecture. However, these methods often require multiple forward passes or model instances, increasing computational and memory costs. This limits their practicality in real-time or resource-constrained environments. Despite these trade-offs, they provide robust calibration and uncertainty quantification when efficiency is not the primary constraint.
@@ 4.4 Hybrid Calibration Methods
Beside applying each method independently, we can always have calibration compositions by combining two or more methods. One straightforward way to combine non-post-hoc methods with post-hoc methods. For instance, performing Temperature Scaling (TS) after employing the regularization method and implicit calibration (?, ?). Thulasidasan et al. (?) observed that the combination of label smoothing and mixup training significantly improved calibration. While there are several possibilities for combining different approaches, we highlight some interesting compositions.
@@ 4.4.1 Ensemble Temperature Scaling (ETS)
Zhang et al., (?) gave three important definitions related to calibration properties: accuracy-preserving, data-efficient, and expressive. They pointed out that TS is an accuracy-preserving and data-efficient approach but is less expressive. Ensemble Temperature Scaling (ETS) was proposed to improve TS expressivity while maintaining the other two properties:
@@ 4.4.2 Temperature Scaling with MC Dropout
[…]
@@
4.4.3 Conformal Prediction with MC Dropout
Cocheteux et al. (?) propose an uncertainty-aware online extrinsic calibration pipeline that integrates MC Dropout for sampling epistemic uncertainty over calibration parameters with Conformal Prediction (?, ?, ?) to generate prediction intervals with statistically guaranteed coverage. They apply MC‑Dropout inference on backbone calibration models, compute nonconformity scores from the dropout-generated samples, and use conformal quantiles to construct confidence intervals. Evaluations on multiple datasets demonstrate that the method achieves near-nominal coverage (PICP), acceptable interval widths (MPIW), and improved interval scores (IS), all while supporting dynamic, online calibration scenarios.
4.4.4 Practical Considerations
Combining different types of calibration methods—particularly post-hoc techniques with regularization-based or uncertainty-aware methods—can yield more robust and flexible calibration pipelines. Such hybrid strategies often leverage complementary strengths: for instance, the simplicity and data efficiency of post-hoc methods like Temperature Scaling (TS) alongside the expressiveness or generalization benefits of training-time techniques. However, these benefits may come at the cost of increased model complexity, interpretability challenges, and higher computational overhead. For example, Ensemble Temperature Scaling (ETS) (?) enhances expressivity but introduces additional parameters and components that must be tuned. Similarly, integrating Monte Carlo (MC) Dropout into post-hoc calibration (as in TS or Conformal Prediction settings) improves uncertainty estimation but requires multiple stochastic forward passes, potentially increasing inference time. Thus, while hybrid methods present promising improvements in calibration quality, their design should carefully consider trade-offs between performance, scalability, and deployment constraints.
@@ 4.5.1 Summary of Strengths and Limitations
Implementation and Tuning Complexity. Implementing regularization method is less straightforward as compared to post-hoc ones. Also this line of work usually introduce hyper-parameters, for example, smoothing factor in label smoothing (?) and  (?) empirically showed that the importance of $\gamma$ in calibrating deep models. However, selecting the appropriate hyper-parameters can be challenging and may require extensive hyper-parameter tuning.
Uncertainty Estimation (UE) (1) Uncertainty quantification. Bayesian-method based calibration can quantify the uncertainty in the calibrated model (?). (2) Incorporating prior information. UE methods allow incorporation of prior knowledge or beliefs about the parameters being estimated and provide more robust estimates, especially when dealing with limited data (?). Computational complexity. Many uncertainty estimation methods involve computationally intensive procedures, which can be impractical for large datasets or real-time applications, limiting their scalability. For example MC-dropout (?) requires multiple inference runs and perform averaging on predictions.
Hybrid Methods Combines strengths: Hybrid methods can leverage the advantages of multiple approaches, potentially leading to better overall performance by addressing the limitations of individual techniques. Section 4.4 describes some of these methods Increased complexity: Hybrid methods usually introduce additional complexity in model calibration, Integrating multiple techniques into a cohesive framework may pose implementation challenges and increase the risk of errors requiring more extensive validation and testing to ensure robustness and effectiveness.
@@ 4.5.2 Backbone Models and Datasets
Table 2: The calibration methods and their categorization, datasets and backbone models to be calibrated with those methods.
5 Calibrating Pre-trained Large Models
@@
5.1 Large Vision Models (LVMs)
Minderer et al. (?) studied recent state-of-the-art vision models that include vision transformer (?) and MLP-mixer (?). They found out that the model size and amount of pre-training in the recent model generation could not fully explain the observed decay of calibration with distribution shift or model size in the prior model generation. They also discussed the correlation between in-distribution and out-of-distribution (OOD) calibration. They pointed out that the models with better in-distribution calibration also gave better calibration on OOD benchmarks. LeVine et al. (?) studied the calibration of CLIP (?) as a zero-shot inference model and found that CLIP is miscalibrated in zero-shot settings.They showed effectiveness of learning a temperature on an auxiliary task and applying it to inference regardless of prompt or datasets.
5.2 Large Language Models (LLMs)
Pre-trained LLMs like GPT-3 and its successors exhibit strong generalization across diverse tasks through in-context learning (ICL). However, their confidence estimates are often poorly aligned with actual correctness, especially in few-shot and zero-shot settings. Unlike traditional supervised models, LLMs are typically used in prompt-based or API-only scenarios where internal logits are not directly accessible, making black-box calibration a key research challenge. This has motivated a growing body of work aimed at reducing overconfidence, mitigating bias, and improving the reliability of LLM predictions across task types and domains.
The outcome of prompting-based learning can be unstable and introduce biases through prompt format and example selection (?). The introduced biases include majority label bias, recency bias, and token bias. To mitigate this, (?) proposed contextual calibration, which improves the predictive power of GPT-3 in few-shot settings by estimating bias using a context-free input such as ”N/A”, and then applying vector scaling (?):
@@ 5.2 Large Language Models (LLMs)
Domain-Context Calibration (DC)  (?) was proposed to mitigate label biases in ICL. Instead of a fixed token like ”N/A”, the method uses random in-domain text to estimate label-wise priors:
@@
Zheng et al. (?) investigated selection bias in multiple-choice tasks and identified token bias as a core factor. Their proposed method, PriDe, is a label-free, inference-time debiasing method with strong efficiency and cross-domain generalization.
Park et al. (?) extended mixup (?) training for pre-trained LMs by generating synthetic examples based on the Area Under the Margin (AUM), improving calibration in classification settings. Another approach called Prototypical calibration (PROCA) (?) emphasizes the importance of decision boundaries. It fits a Gaussian mixture model (GMM) over model outputs:
$P_{GMM}(\hat{p})=\sum_{k=1}^{K}\alpha_{k}p_{G}(\hat{p}|u_{k},\Sigma_{k}),~~~p_{calibrated}=\frac{1}{T}\sum_{t=1}^{T}P_{GMM}(\hat{p}_{t})$ (38)
where parameters $\{\alpha_{k},u_{k},\Sigma_{k}\}_{k=1}^{K}$ are estimated via Expectation-Maximization (?).
Zhou et al. (?) identified limitations in earlier methods. For example, contextual and domain-context calibration struggle with multi-sentence inputs, and PROCA may overfit. They proposed Batch Calibration, which uses class-wise prior estimation over batches to correct contextual bias. Unlike prior techniques that rely on handcrafted calibration prompts or context-free inputs (e.g., ”N/A”), Batch Calibration estimates contextual bias directly from a batch of real task inputs. Specifically, it marginalizes model outputs across a batch to estimate the background distribution over labels, then debiases each individual prediction accordingly:
$\hat{p}(y)=\frac{1}{B}\sum_{i=1}^{B}p(y|x_{i},\textsc{Context}),\quad p_{\text{calibrated}}(y|x)=p(y|x,\textsc{Context})-\hat{p}(y)$ (39)
Here, $B$ denotes the batch size. The method is robust to prompt variations. It unifies previous approaches such as Contextual Calibration and Domain Calibration under a more stable framework, and achieves state-of-the-art calibration and accuracy across multiple tasks using models like PaLM-2 and CLIP.
Shen et al. (?) introduced Thermometer, an auxiliary‑model‑based, post‑hoc calibration method designed specifically for open‑form large language models (LLMs). Unlike common approaches that rely on multiple forward passes or fine‑tuning at inference, Thermometer trains a lightweight, task-agnostic calibration network across multiple datasets. At inference time, this auxiliary model adjusts the LLM’s predicted probabilities to improve ECE without altering the LLM’s outputs or requiring additional expensive inferences. Empirical results on multiple benchmarks show that Thermometer consistently achieves lower calibration error than temperature scaling or MC‑based methods, preserves accuracy, and provides faster inference—thanks to its single-pass inference overhead.
Zhang et al. (?) proposed atomic calibration, a novel framework that decomposes long-form answers into individual, self-contained factual units known as atomic claims, and then estimates a separate confidence score for each. Two types of methods are introduced: generative methods that assess confidence based on sampling consistency across multiple generations, and discriminative methods that query the model directly for confidence in each claim. The authors also introduce new evaluation metrics such as Unnormalized and Quantized Claim Calibration Errors (UCCE and QCCE) to measure calibration quality at both the claim and response levels. Experiments conducted across several open-source LLMs and long-form QA datasets show that atomic calibration yields significantly more accurate and interpretable confidence estimates. Furthermore, combining generative and discriminative signals produces the best overall calibration performance. This work demonstrates the necessity of fine-grained calibration techniques for open-form LLM outputs and provides a scalable framework to achieve it.
6 Calibration Applications
@@
6.1 Medical Imaging and Analysis
Calibrated deep learning models play a crucial role in medical imaging tasks such as disease diagnosis and medical image analysis (?). These models provide accurate predictions and uncertainty estimates, assisting clinicians in making informed decisions about patient care. For instance, calibrated models improve the reliability of automated diagnoses in radiology, pathology, and other medical imaging domains. By aligning predicted probabilities with observed outcomes, these models enhance the trustworthiness of automated medical decision support systems. Numerous studies have demonstrated the effectiveness of model calibration techniques in improving the accuracy and interpretability of medical imaging models. Penso et al.  (?) presents a network calibration procedure to make be robust to label noise based on observation that the confusion matrix of the noisy labels can be expressed as the matrix product between the confusion matrix of the clean labels and the label noises. Ricci Lara et al. (?) identified the issue of discrimination and calibration biases in models trained for automatic detection of malignant dermatological conditions from skin lesions images. They found that calibration metrics may appear to show that model outputs do not fairly represent uncertainty.
6.2 Financial Decision Making